# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

### v3

@@ 4.5 Prototype Representation Fusion Module
Currently, there are two learned hidden representations, one is $\bm{z}$ obtained through the missing-aware self-attention module, and the other is $\bm{\mathcal{G}}$ obtained through the prototype similar patient cohort discovery module. A fusion module thus is needed to comprehensively fuse the information:
@@
Table 1: In-hospital mortality and readmission prediction results on MIMIC-III and MIMIC-IV datasets. The reported score is in the form of mean±std of 5 runs. Bold indicates the top performance. All metrics are multiplied by 100 for readability purposes.
Dataset MIMIC-III Mortality MIMIC-III Readmission MIMIC-IV Mortality MIMIC-IV Readmission
Metric AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$)
MICE 52.40±0.57 85.43±0.23 50.61±0.47 77.97±0.60 49.89±0.78 84.37±0.13 45.28±0.38 75.84±0.25
GRU-D 45.31±3.22 81.72±2.46 42.13±3.19 73.31±1.65 48.79±1.57 85.02±0.50 45.03±0.71 76.47±0.53
RETAIN 51.76±0.86 85.57±0.43 47.53±0.48 77.42±0.38 54.06±0.71 86.24±0.36 49.93±0.73 78.54±0.38
AdaCare 52.28±0.50 85.73±0.19 48.76±0.35 77.65±0.32 50.45±0.80 83.96±0.13 48.57±0.29 77.00±0.20
ConCare 51.45±0.76 86.18±0.14 47.45±0.96 77.74±0.32 49.97±1.08 85.41±0.40 47.17±0.84 77.47±0.19
GRASP 53.59±0.33 86.54±0.17 50.21±0.22 78.14±0.35 54.41±0.46 86.08±0.17 50.22±0.26 78.50±0.22
M3Care 51.68±1.03 86.23±0.42 49.00±0.71 78.00±0.55 52.95±0.71 84.90±0.37 49.22±0.69 77.31±0.42
SAFARI 45.92±1.01 85.10±0.24 45.59±0.35 77.01±0.21 46.58±0.55 46.58±0.55 44.78±0.69 76.05±0.38
AICare 51.37±0.70 85.40±0.48 47.06±1.16 76.23±0.84 49.76±0.86 84.62±0.28 45.88±1.12 76.07±0.43
PRISM-${}_{\text{-proto.}}$ 55.52±0.34 87.28±0.11 51.17±0.25 78.66±0.22 55.76±0.90 86.82±0.16 50.75±0.65 79.12±0.44
PRISM-${}_{\text{-calib.}}$ 56.16±0.42 87.33±0.22 49.13±2.11 77.87±0.88 55.18±0.77 86.57±0.20 50.62±0.70 78.66±0.45
PRISM 57.02±0.38 87.34±0.22 51.31±1.02 78.76±0.59 55.92±0.75 86.82±0.20 51.04±0.70 79.14±0.33
5 Experiments
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are open-sourced for reproducibility.
5.1 EHR Datasets
Both MIMIC-III and MIMIC-IV datasets are split into 70% training set, 10% validation set and 20% test set with stratified shuffle split strategy based on patients’ end-stage mortality outcome. By default, both datasets employ the Last Observation Carried Forward (LOCF) imputation method Woolley et al. (2009).
MIMIC-III Johnson et al. (2016) (Medical Information Mart for Intensive Care) is a large, freely-available database comprising information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality. The subsequent iteration, MIMIC-IV Johnson et al. (2023), stands as an evolved manifestation of the MIMIC-III database, encompassing data updates and partial table reconstructions. We extracted patient EHR data following Harutyunyan et al. (2019).
@@ 5.2 Evaluation Metrics
We assess the binary classification performance using AUROC, AUPRC and F1. Here we emphasize AUPRC as the main metric due to it is informative when dealing with highly imbalanced and skewed datasets Kim and Hwang (2022); Davis and Goadrich (2006) as MIMIC datasets.
@@ 5.3 Baseline Models
include MICE Van Buuren and Groothuis-Oudshoorn (2011) addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model Hochreiter and Schmidhuber (1997). GRU-D Che et al. (2018) incoporates both the last observed and global mean values in the GRU network to handle missing values,
@@
such as RETAIN Choi et al. (2016), AdaCare Ma et al. (2020a), ConCare Ma et al. (2020b), GRASP Zhang et al. (2021), M3Care Zhang et al. (2022), SAFARI Ma et al. (2022) and AICare Ma et al. (2023). They focus on personalized health status embeddings.
@@
include PRISM-${}_{\text{-proto.}}$ that removes the confidence-aware prototype similar patient learner, and PRISM-${}_{\text{-calib.}}$ that removes the feature confidence learner for missing-feature-aware calibration process.
@@ 5.4 Experimental Results
Table 1 depicts the performance evaluation of baseline methods, PRISM, and its reduced versions for ablation study on both the MIMIC-III and MIMIC-IV datasets under two prediction tasks. Additionally, we conduct t-test based on the AUPRC metric, the PRISM’s performance improvement against all models are all statistically significant with p-value $<$ 0.01, which underscores that PRISM significantly outperforms all baseline models. Specifically, PRISM outperforms models focused solely on enhancing feature representations with attention mechanisms (like RETAIN, AdaCare, ConCare, AICare, SAFARI), by integrating missing feature status into these mechanisms for improved attention calibration and feature representation. It also exceeds models using similar patient information (such as GRASP, M3Care), showing the value of missing feature status in refining prototype patient representations for better performance. PRISM’s advantage over GRU-D, which only considers local patient visit-based feature missing status, highlights the significance of a global perspective on overall feature missing rates for effective feature representation across patients.
@@ 5.5 Ablation Study
Compared to reduced versions, PRISM outperforms PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$ on main metric AUPRC. This indicates that the two designed learners can enhance patient feature representations from different perspectives: the patient’s individual health data utilized by the feature confidence learner based on the attention mechanism and the prototype similar patient representations utilized by the prototype similar patient learner.
@@
To deeply explore the impact of components within each module, we conduct experiments in Table 2, showing PRISM outperforms all baselines. The symbol $\phi$ denotes similarity measure, detailed in Equation 14. The term $z$ alone indicates the use of L2 distance for patient similarity measure, whereas $z,C$ additionally incorporates feature confidence, enhancing the model’s discriminative capability. Comparing the roles of various components within the feature confidence learner, the performance when considering both global feature missing rate $\rho$ and local patient’s time interval $\tau$ is higher than considering any single component, which illustrates the necessity of considering the feature missing status from both global and local perspectives. When only considering the local perspective, its performance actually worsens, which is consistent with our observation of the performance of GRU-D. As for confidence-aware prototype patient learner, the performance of confidence-aware patient similarity measurement surpasses that without considering feature confidence, which also shows the impact of missing feature status on measuring similar patients.
Table 2: Performance comparison of internal components on the MIMIC-IV mortality prediction task. The reported score is in the form of mean±std. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the highest performance within each components, Red denotes the highest performance among all comparisons.
@@ 5.6 Observations and Analysis
Figure 2: AUPRC Performance across 5 Sparsity Levels. PRISM significantly outperforms other models in extreme sparse scenario.
@@
To assess PRISM’s performance in conditions of data scarcity, we compare it with leading models such as GRASP, MICE(LSTM), and RETAIN in the MIMIC-IV mortality prediction task. We intentionally increase the missing data rates beyond the original missing rate in MIMIC-IV. The results in Figure 2 demonstrate that PRISM excels across all conditions. Notably, in situations of extreme data scarcity, PRISM significantly outperforms the other models, highlighting its robustness in sparse data.
@@
We conduct a detailed analysis to examine the impact of prototype patients cohort diversity and the role of a missing-feature-aware module in patient prototypes. Figure 3 shows that integrating a missing-feature-aware module consistently enhances performance across various cluster sizes, as indicated by the superior AUPRC and F1 scores. Furthermore, the relatively consistent performance across different cluster sizes demonstrates that our model is not overly sensitive to the number of clusters, highlighting its adaptability and robustness in managing various cohort sizes.
@@
Figure 3: Prototype patient cohort size (MIMIC-IV, mortality). With the missing-feature-aware module, PRISM outperforms its counterpart without. It also shows PRISM is not sensitive to the cohort size.
@@
Table 3: Different similarity measures (MIMIC-IV, mortality).
@@
In our comparative study, we evaluate the efficiency and complexity against SOTA EHR-specific models baseline models. Table 4 highlights PRISM’s balance between a relatively low total parameter count (0.882M) and a competitive runtime (69.03s).
Table 4: Computational comparison with latest SOTA EHR-specific models on MIMIC-IV. All models employ the same 128 hidden dimension and 1024 batch size hyperparameters.
@@
ConCare (AAAI-2020) 13.019M 96.62s
GRASP (AAAI-2021) 13.153M 97.68s
M3Care (KDD-2022) 1.351M 68.29s
SAFARI (TKDE-2022) 7.228M 78.97s
AICare (Patterns-2023) 8.901M 72.97s
PRISM 0.882M 69.03s
@@
Cross-Feature Attention Map
Figure 4 presents the cross-feature attention map from PRISM, contrasting average attention weights with and without feature confidence calibration process. This visualization, based on a single diagnostic record from randomly selected MIMIC-IV patients, plots Key features on the x-axis against Query features on the y-axis. Notably, PRISM reduces attention on features like capillary refill rate, fraction inspired oxygen, and height, which have high missing rates (99.64%, 93.62%, and 99.61%, respectively) in the dataset. PRISM recognizes this characteristic, thereby causing the three horizontal lines on the right side of the graph to appear distinctively whiter.

### v4

@@ 4.5 Prototype Representation Fusion Module
Currently, there are two learned hidden representations, one is $\bm{z}$ obtained through the missing-aware self-attention module, and the other is $\bm{\mathcal{G}}$ obtained through the prototype similar patient cohort discovery module. Thus, the patient representation is fused as:
@@
5 Experimental Setups
5.1 Benchmarked Real-World Datasets
We employ MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets datasets for benchmarking. All 4 datasets are split into 70% training set, 10% validation set and 20% test set with stratified shuffle split strategy based on patients’ end-stage mortality outcome. By default, we use the Last Observation Carried Forward (LOCF) imputation method [21].
  1. 1.
    MIMIC-III [22] (Medical Information Mart for Intensive Care) is a large, freely-available database comprising information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality.
  2. 2.
    MIMIC-IV [23] as MIMIC-III dataset’s subsequent iteration, stands as an evolved manifestation of the MIMIC-III database, encompassing data updates and partial table reconstructions. We extracted patient EHR data following [24].
  3. 3.
    PhysioNet Challenge 2012 [25] (Challenge-2012) covers records from 12,000 adult ICU stays. The challenge is designed to promote the development of effective algorithms for predicting in-hospital mortality based on data from the first 48 hours of ICU admission. We utilize 35 lab test features and 5 demographic features in Challenge-2012 dataset.
  4. 4.
    eICU [26] is a large-scale, multi-center ICU database derived from over 200,000 ICU admissions across the United States between 2014 and 2015. 12 lab test features and 2 demographic features are adopted.
The statistics of datasets is in Table 2.
Table 2: Statistics of datasets after preprocessing. The proportion demonstrates the percentage of the label with value $1$. $Out.$ denotes Mortality Outcome, $Re.$ denotes Readmission.
Dataset MIMIC-III MIMIC-IV Challenge-2012 eICU
# Samples 41517 56888 4000 73386
Missing 69.87% 74.70% 84.68% 42.61%
$\text{Label}_{Out.}$ 10.62% 9.55% 13.85% 8.32%
$\text{Label}_{Re.}$ 14.74% 13.85% / /
@@ 5.2 Evaluation Metrics
We assess the binary classification performance using AUROC, AUPRC and F1. Here we emphasize AUPRC as the main metric due to it is informative when dealing with highly imbalanced and skewed datasets [27, 28] as shown in our selected datasets.
@@ 5.3 Baseline Models
We include imputation-based methods, EHR-specific models, and PRISM model with reduced modules as baseline models.
@@
include MICE [12] addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model [29]. GRU-D [10] incoporates both the last observed and global mean values in the GRU network. Additionally, GRU-D utilizes an exponential decay mechanism to manage the temporal dynamics of missing values.
@@
such as RETAIN [30], AdaCare [15], ConCare [1], GRASP [2], M3Care [11], SAFARI [4] and AICare [16]. They focus on personalized health status embeddings.
@@
include PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$. Specifically, PRISM-${}_{\text{-proto.}}$ removes the confidence-aware prototype similar patient learner and reserves the feature-missing-aware calibration process. We apply the patients hidden representation $\bm{z}$ for downstream GRU module. PRISM-${}_{\text{-calib.}}$ removes the feature confidence learner from the missing-feature-aware calibration process. As the confidence-aware prototype patient learner requires the feature confidence $\bm{C}$ as input, we set the $\bm{C}=1$ for each feature.
5.4 Implementation Details
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s RAM size is 64GB. We implement the model in Python 3.11.4, PyTorch 2.0.1 [31], PyTorch Lightning 2.0.5 [32], and pyehr [33]. AdamW [34] is employed with a batch size of 1024 patients. All models are trained for 50 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers. Performance is reported in the form of mean±std of 5 runs with random seeds ${0,1,2,3,4}$ for MIMIC-III and MIMIC-IV datasets, and apply bootstrapping on all test set samples 10 times for the Challenge-2012 and eICU datasets, following practices in [16]. All metric scores are multiplied by 100 for readability purposes.
6 Experimental Results and Analysis
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets, in-hospital mortality prediction task on Challenge-2012 and eICU datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are open-sourced for reproducibility.
Table 3: Benchmarking results on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets. Bold indicates the best performance.
Dataset MIMIC-III Mortality MIMIC-III Readmission MIMIC-IV Mortality MIMIC-IV Readmission Challenge-2012 Mortality eICU Mortality
Metric AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$) AUPRC($\uparrow$) AUROC($\uparrow$)
MICE 52.40±0.57 85.43±0.23 50.61±0.47 77.97±0.60 49.89±0.78 84.37±0.13 75.84±0.25 45.28±0.38 32.26±2.85 72.41±1.46 44.44±4.16 85.19±2.97
GRU-D 45.31±3.22 81.72±2.46 42.13±3.19 73.31±1.65 48.79±1.57 85.02±0.50 76.47±0.53 45.03±0.71 25.43±3.27 63.75±1.76 42.59±3.33 83.85±2.31
RETAIN 51.76±0.86 85.57±0.43 47.53±0.48 77.42±0.38 54.06±0.71 86.24±0.36 78.54±0.38 49.93±0.73 30.23±2.24 69.82±1.89 39.89±3.08 82.53±2.45
AdaCare 52.28±0.50 85.73±0.19 48.76±0.35 77.65±0.32 50.45±0.80 83.96±0.13 77.00±0.20 48.57±0.29 33.10±3.42 69.66±1.57 42.48±3.61 83.91±2.32
ConCare 51.45±0.76 86.18±0.14 47.45±0.96 77.74±0.32 49.97±1.08 85.41±0.40 77.47±0.19 47.17±0.84 30.24±3.00 70.19±2.54 44.40±4.35 85.05±2.93
GRASP 53.59±0.33 86.54±0.17 50.21±0.22 78.14±0.35 54.41±0.46 86.08±0.17 78.50±0.22 50.22±0.26 26.03±3.43 67.14±2.61 45.41±4.05 85.69±2.38
M3Care 51.68±1.03 86.23±0.42 49.00±0.71 78.00±0.55 52.95±0.71 84.90±0.37 77.31±0.42 49.22±0.69 32.63±2.54 73.26±1.67 44.95±4.32 85.44±2.56
SAFARI 45.92±1.01 85.10±0.24 45.59±0.35 77.01±0.21 46.58±0.55 46.58±0.55 76.05±0.38 44.78±0.69 28.94±3.16 70.67±1.76 35.26±3.59 80.10±2.40
AICare 51.37±0.70 85.40±0.48 47.06±1.16 76.23±0.84 49.76±0.86 84.62±0.28 76.07±0.43 45.88±1.12 23.99±2.48 67.35±2.20 42.80±3.79 84.26±2.64
PRISM-${}_{\text{-proto.}}$ 55.52±0.34 87.28±0.11 51.17±0.25 78.66±0.22 55.76±0.90 86.82±0.16 79.12±0.44 50.75±0.65 30.92±2.95 68.87±2.61 50.03±3.97 85.93±1.67
PRISM-${}_{\text{-calib.}}$ 56.16±0.42 87.33±0.22 49.13±2.11 77.87±0.88 55.18±0.77 86.57±0.20 78.66±0.45 50.62±0.70 30.42±2.96 71.07±2.13 46.92±3.20 84.69±1.39
PRISM 57.02±0.38 87.34±0.22 51.31±1.02 78.76±0.59 55.92±0.75 86.82±0.20 79.14±0.33 51.04±0.70 33.60±3.41 73.47±1.11 50.41±3.63 85.82±1.43
@@ 6.1 Experimental Results
Table 3 depicts the performance evaluation of baseline methods, PRISM, and its reduced versions for ablation study on four datasets under two prediction tasks. Additionally, we conduct t-test based on the AUPRC metric, the PRISM’s performance improvement against all models are all statistically significant with p-value $<$ 0.01, which underscores that PRISM significantly outperforms all baseline models. Specifically, PRISM outperforms models focused solely on enhancing feature representations with attention mechanisms (like RETAIN, AdaCare, ConCare, AICare, SAFARI), by integrating missing feature status into these mechanisms for improved attention calibration and feature representation. It also exceeds models using similar patient information (such as GRASP, M3Care), showing the value of missing feature status in refining prototype patient representations for better performance. PRISM’s advantage over GRU-D, which only considers local patient visit-based feature missing status, highlights the significance of a global perspective on overall feature missing rates for effective feature representation across patients.
@@ 6.2 Ablation Study
PRISM outperforms PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$ on main metric AUPRC. This indicates that the two designed learners can enhance patient feature representations from different perspectives: the patient’s individual health data utilized by the feature confidence learner based on the attention mechanism and the prototype similar patient representations utilized by the prototype similar patient learner.
@@
To deeply explore the impact of components within each module, we conduct experiments in Table 4, showing PRISM outperforms all baselines. The symbol $\phi$ denotes similarity measure, detailed in Equation 14. The term $\bm{z}$ alone indicates the use of L2 distance for patient similarity measure, whereas $\bm{z},\bm{C}$ additionally incorporates feature confidence, enhancing the model’s discriminative capability. Comparing the roles of various components within the feature confidence learner, the performance when considering both global feature missing rate $\rho$ and local patient’s time interval $\tau$ is higher than considering any single component, which illustrates the necessity of considering the feature missing status from both global and local perspectives. When only considering the local perspective, its performance actually worsens, which is consistent with our observation of the performance of GRU-D. As for confidence-aware prototype patient learner, the performance of confidence-aware patient similarity measurement surpasses that without considering feature confidence, which also shows the impact of missing feature status on measuring similar patients.
Table 4: Performance comparison of internal components on the MIMIC-IV mortality prediction task. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the best performance within each components, Red denotes the highest performance among all comparisons.
@@ 6.3 Observations and Analysis
To assess PRISM’s performance under conditions of data sparsity, we compare it with leading models such as GRASP, MICE(LSTM), and RETAIN in MIMIC-IV in-hospital mortality prediction task. At the volume-level in Figure 2(a), we reduce the data samples in training set, while at the feature-level in Figure 2(b), we intentionally increase the missing feature rates beyond the original missing rate. PRISM excels in both settings. Notably, in situations of extreme data sparsity, such as using only 10% training data and with an overall 97.47% feature missing rate, PRISM significantly outperforms the other models, highlighting its robustness in handling sparse data.
(a) Volume-level Sparsity
(b) Feature-level Sparsity
Figure 2: AUPRC performance across 5 sparsity levels in MIMIC-IV in-hospital mortality prediction task. PRISM significantly outperforms other models in extremely sparse scenarios on both sparsity settings.
@@
We conduct a detailed analysis to examine the impact of prototype patients cohort diversity and the role of a missing-feature-aware module in patient prototypes. Figure 3 shows that integrating a missing-feature-aware module consistently enhances performance across various cluster sizes, as indicated by the superior AUPRC and F1 score. Furthermore, the relatively consistent performance across different cluster sizes demonstrates that our model is not overly sensitive to the number of clusters, highlighting its adaptability and robustness in managing various cohort sizes.
@@
Figure 3: AUPRC and F1 score performance on various prototype patient cohort size in MIMIC-IV in-hospital mortality prediction task. With the missing-feature-aware module, PRISM outperforms its counterpart without. It also shows PRISM is not sensitive to the cohort size.
@@
Table 5: Comparing different similarity measures in MIMIC-IV in-hospital mortality prediction task.
@@
We evaluate the efficiency and complexity against SOTA EHR-specific models baseline models. Table 6 highlights PRISM’s balance between a relatively low total parameter count (215K) and a competitive runtime (69.03s).
Table 6: Computational comparison with latest SOTA EHR-specific models on the MIMIC-IV dataset. All models employ the same 128 hidden dimension and 1024 batch size hyperparameters.
@@
ConCare (AAAI-2020) 3.3M 96.62s
GRASP (AAAI-2021) 3.3M 97.68s
M3Care (KDD-2022) 469K 68.29s
SAFARI (TKDE-2022) 1.8M 78.97s
AICare (Patterns-2023) 2.2M 72.97s
PRISM 215K 69.03s
Table 7 showcase PRISM’s data preparation time in preprocessing and the overall time for training&validation, and testing on MIMIC datasets. The PRISM pipeline seamlessly computes feature missing statuses in the LOCF pipeline with little extra cost. Moreover, the overall pipeline can be completed in 10 minutes.
Table 7: Data preparation time comparison. Prep. denotes Preprocessing, Val. denotes Validation. All with 50 epochs of training, validate at the end of each epoch. The units are seconds.
Dataset MIMIC-III MIMIC-IV
Prep. w/o Impute LOCF +Ours w/o Impute LOCF +Ours
Time 77.06 300.63 432.19 104.67 358.29 502.68
Pipeline Data Prep. Train+Val. Test Data Prep. Train+Val. Test
Time 432.19 505.83 10.25 502.68 510.73 12.39
Cross-Feature Attention Map
Figure 4 presents the cross-feature attention map from PRISM, contrasting average attention weights with and without feature confidence calibration process. This visualization, based on a single diagnostic record from randomly selected MIMIC-IV patients, plots Key features on the x-axis against Query features on the y-axis. Notably, PRISM reduces attention on features like capillary refill rate, fraction inspired oxygen, and height, which have high missing rates (99.64%, 93.62%, and 99.61%, respectively) in the dataset. PRISM accurately recognizes and calibrates the features with high missing rates, thereby causing the three horizontal lines on the right side of the graph to appear distinctively whiter, showcasing attention-based feature learner module’s interpretability which other baselines lack.
@@
Feature Decay Rates Observation
Figure 5 displays the decay rates of adaptive learning for various features, indicating how their importance diminishes over time. Higher decay rates suggest that the model prioritizes immediate changes in features like heart rate and pH, which is crucial for detecting acute medical conditions such as shock or infection. Conversely, features like sex and systolic blood pressure exhibit lower decay rates, highlighting their relevance in long-term analysis. Notice that the feature ‘Height’ exhibits short-term dynamics, likely due to being absent in over 99% of the data. Consequently, the rapid decay of the ‘Height’ feature, resulting from its high missing rate, does not significantly influence clinical decisions over the long term, which aligns with our intuition.
Figure 5: Feature decay rates observation